# PrefixKV: Adaptive Prefix KV Cache is What Vision Instruction-Following Models Need for Efficient Generation

### v3

@@ Abstract
Recently, large vision-language models (LVLMs) have rapidly gained popularity for their strong generation and reasoning capabilities given diverse multimodal inputs. However, these models incur significant computational and memory overhead during inference, which greatly hinders the efficient deployment in practical scenarios. The extensive key-value (KV) cache, necessitated by the lengthy input and output sequences, notably contributes to the high inference cost. Based on this, recent works have investigated ways to reduce the KV cache size for higher efficiency. Although effective, they generally overlook the distinct importance distributions of KV vectors across layers and maintain the same cache size for each layer during the next token prediction. This results in the significant contextual information loss for certain layers, leading to notable performance decline. To address this, we present PrefixKV. It reframes the challenge of determining KV cache sizes for all layers into the task of searching for the optimal global prefix configuration. With an adaptive layer-wise KV retention recipe based on binary search, the maximum contextual information can thus be preserved in each layer, facilitating the generation. Extensive experiments demonstrate that our method achieves the state-of-the-art performance compared with others. It exhibits superior inference efficiency and generation quality trade-offs, showing promising potential for practical applications. Code is available at https://github.com/THU-MIG/PrefixKV.
@@ 1 Introduction
Recent years have witnessed the significant advancements of large vision-language models (LVLMs) [37, 72, 26, 36, 2, 15, 11]. Based on the powerful Large Language Models (LLMs) [18, 46, 51, 1, 50, 28, 3, 63], these models integrate visual inputs, showing strong generation and reasoning abilities for various multimodal tasks. They demonstrate inspiring application potential in various fields like autonomous driving [14, 57] and intelligent medical analyses [29, 52].
Figure 1: Comparison between previous methods and ours. Previous methods simply keep the same prefix length for priority sequences of KV, i.e., retraining the same cache size for each layer. This causes notable information loss for certain layers. In this example, the first layer loses 30% of information. In contrast, we derive the optimal global prefix configuration to preserve as much information as possible in each layer. In this example, both layers can retain 90% of information, thereby enhancing performance.
@@
However, despite their remarkable capabilities, the efficient deployment of LVLMs encounters notable challenges in real-world scenarios. This stems from the typical Transformer architecture employed in the LVLMs, which necessitates the global interaction of tokens. During autoregressive decoding, the key and value vectors of previous tokens are thus required to be stored as the KV cache and subsequently retrieved by the output token [42]. The KV cache grows linearly with the number of processed tokens, which leads to notable memory overhead and heavy burden on GPU communication under lengthy sequences. This causes suboptimal efficiency, resulting in the inference bottleneck.
@@
Given this, recent works have investigated pruning unimportant KV vectors or merging adjacent vectors to reduce the KV cache size while preserving the model performance [39, 75, 61]. For example, H2[75] discards less important ones based on the attention scores. Elastic Cache [39] identifies the important KV vectors as the anchor points and merges the surrounding less important cache with these anchors. While effective, existing works typically apply a uniform strategy that retains the same number of KV vectors for each layer to generate the next token efficiently, overlooking the layer-wise heterogeneous characteristics. Our analyses in Section 3.2 reveal the notably distinct importance distributions of KV vectors across layers, highlighting the need for tailored retention recipe for each layer. Intuitively, the importance distribution is concentrated in certain layers while relatively dispersed in other layers. As a suboptimal solution, retaining the same cache size for each layer during the next token prediction results in obvious information loss in dispersed layers, and redundant cache in concentrated layers, as shown in Figure 1.(a).
To address this, we present PrefixKV, a new KV cache compression method for efficient and accurate model generation. We define the prefix [^1] Prefix refers to top elements sorted by importance instead of position. of KV cache as the top ones in the priority sequences that are sorted KV vectors according to the normalized importance. The prefix KV vectors of each layer represents the retained cache, making the challenge of determining the optimal layer-wise cache size for the next token generation equivalent to identifying the optimal global prefix configuration. To derive such configuration for a given compression ratio budget, we leverage the prefix cumulative priority as the measurement for the amount of reserved contextual information in each layer. Binary search is then utilized to obtain the desirable information retention ratio, enabling the layer-wise KV retention aligns with the overall budget and maintains the ideal cumulative priority. This ensures that each layer can preserve maximal contextual information after compression, keeping the high generation quality of models, as shown in Figure 1.(b). Extensive experiments show that our method achieves the state-of-the-art performance compared with existing works. It can greatly boost the inference efficiency while well maintaining the model’s strong capabilities, exhibiting promising potential for real-world applications. Notably, with the compression budget of 20%, it provides 1.8$\times$ inference speedup for LLaVA-1.5-7B, attaining the competitive performance compared with original model.
@@ 2 Related Work
2.1 Vision instruction-following model
@@
The progress of large vision-language models (LVLMs) has significantly expanded the capabilities of large language models (LLMs) [1, 50, 51, 18, 28, 2, 47, 48, 46] by incorporating visual information, leading to powerful generation and reasoning ability for multimodal tasks [37, 36, 15, 79, 9, 10, 8, 31, 4]. These models typical use linear projection [37] or perceivers [27] to integrate visual representations into the input of LLMs directly. Then, they are further finetuned on high-quality instructional datasets, which include the image-text pairs and the language instructional commands. These models can thus follow multimodal instructions effectively and accurately. Inspired by the notable advancements, subsequent works have sought to enhance their region-level grounding ability [6, 72, 41, 11], 3D world perception [23, 64, 12, 32], semantic understanding [74, 70, 53, 16], and video comprehension [71, 33, 62, 30], etc. Other works also explore the applications of LVLMs in various practical scenarios, including industrial anomaly recognition [22, 65, 7], biomedical image understanding [29, 52], and autonomous driving and map systems [14, 49, 67], etc.
2.2 KV cache compression
@@
While powerful, existing LVLMs typically encounter high inference costs due to the large parameter count and complex computation, which hinder their deployment for practical applications significantly. Like research on efficient vision models [59, 54, 24, 55, 56], numerous efforts have focused on improving the inference efficiency for LVLMs, including architectural design [13, 77, 69, 80], quantization [35, 60, 38], distillation [25, 21, 43], and pruning [40, 45, 44], etc. These methods typically focus on the compression on the model level, reducing the inference overhead brought by the parameters. Meanwhile, the KV cache on the data level also incurs significant memory usage and GPU communication overhead during inference. To address this, recent works have investigated various ways to compress the KV cache. For example, StreamingLLM [61] leverages the distance to the output token as the importance indicator and preserves the starting KV vectors and those adjacent to the output token. H2[75] utilizes the attention scores as the importance metric for KV vectors and retains the important ones during generation. Elastic Cache [39] divides the KV cache into several buckets according the positions of important KV cache and merges the vectors in the same bucket into one. Despite effective, they often overlook the distinct importance distribution across layers and simply retain the same KV cache size in each layer. This causes the notable information loss and significantly affects the model’s generation capability.
@@ 3 Methodology
In Section 3.1, we first introduce the basic inference process of LVLMs and the general KV cache compression framework. Then, in Section 3.2, we formalize the process of KV compression as retaining the prefix of KV cache and uncover the fact of diverse importance distributions of KV vectors in layers. This variation causes the obvious contextual information loss in existing methods. To address this, we present PrefixKV in Section 3.3, to deliver the ideal layer-wise cache size for the next token prediction by binary searching for the optimal global prefix configuration.
@@ 3.1 Preliminary
The general KV cache compression framework in existing works consist of two stages [39, 75, 61]  [^2] For more preliminary, please refer to previous works [42, 76, 66, 78].. Firstly, after prefilling, the importance of each KV vector is derived based on the attention scores or the distance to the generated output. Then, the most important ones are retained in the cache while the less important KV vectors are removed to reduce the memory footprint. Meanwhile, the reserved KV cache sizes in layers meet the requirement of the compression ratio budget. Secondly, during the decoding stage, with the inclusion of KV vectors in the cache for newly generated tokens, the importance metric of each KV vector is updated. Less critical ones are removed to ensure that the cache size consistently aligns with the overall budget. Our method also follows this general framework, as shown in Figure 3. Specifically, after the prefilling, we retain the most important KV vectors by the ideal global prefix configuration. In the decoding, we follow [39] to prune the vectors at a fixed distance to the latest generated token and maintain the global prefix configuration to meet the target budget.
@@ 3.2 Layer-wise KV Cache Importance Distribution
Mathematical notations for KV compression. We first introduce the necessary notations to formalize the compression procedure. Specifically, we follow [39, 75] to employ the attention scores as the importance metric, which indicates the amount of contextual information of each KV vector. We suppose that the model consists of $L$ transformer layers. For the $l$-th, layer, its input tokens $\{\boldsymbol{t}_{l}^{1},\boldsymbol{t}_{l}^{2},...,\boldsymbol{t}_{l}^{N}\}$ interact with each other in the multi-head self-attention module, where $N$ is the token number. For the $i$-th head, we denote the query, key, and value vectors of the token $\boldsymbol{t}_{l}^{n}$ as $\boldsymbol{q}_{l}^{i,n}$, $\boldsymbol{k}_{l}^{i,n}$, and $\boldsymbol{v}_{l}^{i,n}$, respectively. Then, the causal attention score matrix $\boldsymbol{A}_{l}^{i}=\{\boldsymbol{a}_{l}^{i,m,n}\}_{N\times N}$ can be derived by
$\small\boldsymbol{a}_{l}^{i,m,n}=\frac{\text{exp}(\boldsymbol{q}_{l}^{i,m}\cdot\boldsymbol{k}_{l}^{i,n})}{\sum_{j\leq m}\text{exp}(\boldsymbol{q}_{l}^{i,m}\cdot\boldsymbol{k}_{l}^{i,j})},$ (1)
where $\boldsymbol{a}_{l}^{i,m,n}$ indicates the attention score of token $\boldsymbol{t}_{l}^{m}$ with respect to token $\boldsymbol{t}_{l}^{n}$ in the $i$-th head. Then, the total attention score that each token $\boldsymbol{t}_{l}^{n}$ receives in the $i$-th head can be derived by $\sum_{m}\boldsymbol{a}_{l}^{i,m,n}$. For each $\boldsymbol{k}_{l}^{i,n}$ and $\boldsymbol{v}_{l}^{i,n}$, we follow [39] to define their importance metric as the averaged total attention score of $\boldsymbol{t}_{l}^{n}$ across all heads, i.e.,
$\small\boldsymbol{I}_{l}^{n}=\text{Average}_{i}(\sum_{m}\boldsymbol{a}_{l}^{i,m,n}),$ (2)
where $\text{Average}_{i}$ denotes the average operation for heads. It is noted that the importance metric of each KV vector varies at each layer but is the same across heads. Then, for a compression ratio budget $r$, the top $\boldsymbol{R}_{l}$ proportion of KV vectors with the highest importance are retained in each head at the $l$-th layer. Besides, the adoption of $\{\boldsymbol{R}_{1},\boldsymbol{R}_{2},...,\boldsymbol{R}_{L}\}$ satisfy the requirement of $\sum_{l=1}^{L}\boldsymbol{R}_{l}N=rLN$.
@@
Figure 3: (a) The inference process of LVLMs, where the orange and green rectangles denote the KV cache generated during prefilling and utilized during decoding, respectively. After prefilling, the KV cache is layer-wisely compressed according to the proportions specified by PrefixKV, i.e., $\{\boldsymbol{R}_{1},...,\boldsymbol{R}_{L}\}$. During decoding, as the sequence lengthens and cache increases, the KV cache consistently maintains the derived compression proportions by pruning KV at a fixed distance [39]. (b) The overview of PrefixKV. It employs binary search for cumulative priority sequences of KV to derive the optimal global prefix configuration, which delivers ideal cache size ratio for each layer.
@@
[…]
$\boldsymbol{P}_{l}^{o}$ indicates the amount of contextual information kept in the $l$-th layer after retaining top $o$ proportion of the most important KV vectors. By deriving the prefix size ratio sequence $\{\frac{1}{N},\frac{2}{N},...,1\}$ and its corresponding cumulative priority sequence $\{\boldsymbol{P}_{l}^{\frac{1}{N}},\boldsymbol{P}_{l}^{\frac{2}{N}},...,\boldsymbol{P}_{l}^{1}\}$, we can then obtain the lorenz curve of the importance distribution. As shown in Figure 2, we observe that the cumulative priority growth trends vary significantly across layers. Previous works generally retain the same layer-wise cache size, i.e., adopting the same prefix size ratio of KV cache with $\boldsymbol{R}_{j}=r,\forall j\in[1,L]$. In this scenario, as shown in the dashed black line, different layers exhibit the markedly distinct cumulative priorities, i.e., $\boldsymbol{P}_{l}^{\boldsymbol{R}_{j}}$. This suggests an uneven retention of contextual information across layers, where layers showing rapid growth trends retain a substantial amount, whereas those with slower growth trends retain relatively little. This disparity leads to the obvious information loss in layers with slow growth and adversely impacts the generation quality. We also quantitatively show the diverse importance distribution by the gini coefficient [17, 20]. It is defined as the area that lies between the equality line (the dotted pink line in Figure 2) and the lorenz curve. The smaller it is, the more uniform the importance distribution, and vice versa. As shown in the legend in Figure 2, different layers exhibit varying gini coefficients, further demonstrating the heterogeneity of KV cache in layers. This calls for adaptively determining the prefix of KV cache for each layer.
@@ 3.3 PrefixKV
Based on the above observations, we introduce PrefixKV. With layer-wise cache size scheme formalized as the global prefix configuration, it employs binary search to derive the ideal solution, as shown in Figure 3.(b).
Global prefix configuration. All layers’ cache size ratios $\{\boldsymbol{R}_{1},...,\boldsymbol{R}_{L}\}$, i.e., the prefix size ratios of KV cache, constitute the global prefix configuration space of the model. The goal is to identify the optimal global prefix configuration to maintain the high quality of model generation. Given the compression ratio budget $r$, we discover a information retention ratio $p$ to derive such configuration. Specifically, with the priority sequences, $p$ represents the cumulative priority threshold. For the $l$-th layer, its proportion of retained KV vectors is thus the minimum prefix size ratio $o$ such that the cumulative priority $\boldsymbol{P}_{l}^{o}$ in Equation 3 is larger than or equal to $p$, i.e., $\boldsymbol{R}_{l}=\text{min}(\{o|\boldsymbol{P}_{l}^{o}\geq p\})$. Meanwhile, the value of $p$ satisfies the requirement of the compression ratio budget, i.e.,
@@
1 Initialize $p_{1}\leftarrow 0$, $p_{2}\leftarrow 1$;
2 while $p_{1}<p_{2}$ do
3       $p\leftarrow\frac{p_{1}+p_{2}}{2}$, $\sum_{l}\boldsymbol{R}_{l}=\sum_{l}\text{min}(\{o|\boldsymbol{P}_{l}^{o}\geq p\})$;
4       $\delta=\sum_{l}\boldsymbol{R}_{l}-rL$;
5       if $\delta==0$ then  return $p$ ;
6       else if $\delta<0$ then  $p_{1}\leftarrow p$ ;
7       else  $p_{2}\leftarrow p$ ;
8
9return $p$;
Algorithm 1 Binary Search for retention ratio $p$
Binary search for optimal configuration. Due to the numerous possible values of $p$, obtaining $p$ is quite challenging. We propose employing binary search to efficiently derive $p$ for adaptive layer-wise KV retention. Formally, we start with the interval of $[p_{1},p_{2}]$ where $p_{1}=0$ and $p_{2}=1$. We try $p=\frac{p_{1}+p_{2}}{2}=0.5$ and calculate its corresponding compression budget difference $\delta=\sum_{l}\boldsymbol{R}_{l}-rL$. If $\delta$ is equal to zero, we have found the value for $p$. If it is smaller than zero, we set $p_{1}=p$; otherwise, we set $p_{2}=p$. Then, we update $p=\frac{p_{1}+p_{2}}{2}$ and repeat the process until meeting the constraint. Algorithm 1 presents the process. In this way, as shown in the dashed red line in Figure 2, different layers can consistently keep the maximal amount of the contextual information. It ensures that the valuable KV vectors are cached and accessible during generation across all layers. The quality of the model’s outputs can be maintained even after significant compression of the KV cache. In practice, $p$ that satisfies $\delta=0$ may not exist. We can thus set a small threshold for $\delta$ to terminate the search and scale the resulting global prefix configuration to meet the budget.
Moreover, we observe that the cumulative priority sequences of layers are similar and robust across different samples. Therefore, given a compression ratio budget, we can leverage random samples to derive the corresponding $p$ value and the optimal global prefix configuration $\{\boldsymbol{R}_{1},\boldsymbol{R}_{2},...,\boldsymbol{R}_{L}\}$ for each layer offline, as analyzed in Figure 4, Table 5, and Table 6. The configuration can thus be adopted for the model during inference, which avoids the online binary search and shows good generalizability.

### v4

@@ Abstract
Recently, large vision-language models (LVLMs) have rapidly gained popularity for their strong generation and reasoning capabilities given diverse multimodal inputs. However, these models incur significant computational and memory overhead during inference, which greatly hinders the efficient deployment in practical scenarios. The extensive key-value (KV) cache, necessitated by the lengthy input and output sequences, notably contributes to the high inference cost. Based on this, recent works have investigated ways to reduce the KV cache size for higher efficiency. Although effective, they generally overlook the distinct importance distributions of KV vectors across layers and maintain the same cache size for each layer during the next token prediction. This results in the significant contextual information loss for certain layers, leading to notable performance decline. To address this, we present PrefixKV, where "Prefix" means the top-ranked KV based on importance rather than position in the original sequence. It reframes the challenge of determining KV cache sizes for all layers into the task of searching for the optimal global prefix configuration. With an adaptive layer-wise KV retention recipe based on binary search, the maximum contextual information can thus be preserved in each layer, facilitating the generation. Extensive experiments demonstrate that our method achieves the state-of-the-art performance compared with others. It exhibits superior inference efficiency and generation quality trade-offs, showing promising potential for practical applications. Code is available at https://github.com/THU-MIG/PrefixKV.
@@ 1 Introduction
Recent years have witnessed the significant advancements of large vision-language models (LVLMs) liu2024visual ; zhang2023gpt4roi ; idefics ; liu2024improved ; bai2023qwen ; instructblip ; chen2023shikra . Based on the powerful Large Language Models (LLMs) dubey2024llama ; team2023gemini ; touvron2023llama2 ; achiam2023gpt ; touvron2023llama ; jiang2023mistral ; bai2023qwenllm ; yang2023baichuan , these models integrate visual inputs, showing strong generation and reasoning abilities for various multimodal tasks. They demonstrate inspiring application potential in various fields like autonomous driving cui2024survey ; wang2023drivemlm and intelligent medical analyses li2024llava ; liu2023medical .
@@
However, despite their remarkable capabilities, the efficient deployment of LVLMs encounters notable challenges in real-world scenarios. This stems from the typical Transformer architecture employed in the LVLMs, which necessitates the global interaction of tokens. During autoregressive decoding, the key and value vectors of previous tokens are thus required to be stored as the KV cache and subsequently retrieved by the output token pope2023efficiently . The KV cache grows linearly with the number of processed tokens, which leads to notable memory overhead and heavy burden on GPU communication under lengthy sequences. This causes suboptimal efficiency, resulting in the inference bottleneck.
@@
Given this, recent works have investigated pruning unimportant KV vectors or merging adjacent vectors to reduce the KV cache size while preserving the model performance liu2024efficient ; zhang2024h2o ; xiao2023efficient ; zhang2024pyramidkv . For example, H2zhang2024h2o discards less important ones based on the attention scores. Elastic Cache liu2024efficient identifies the important KV vectors as the anchor points and merges the surrounding less important cache with these anchors. While effective, existing works liu2024efficient ; zhang2024h2o ; xiao2023efficient typically apply a uniform strategy that retains the same number of KV vectors for each layer to generate next token efficiently, often overlooking the layer-wise heterogeneous characteristics. Our analyses in Sec.˜3.2 reveal the notably distinct importance distributions of KV vectors across layers, highlighting the need for tailored retention recipe for each layer adaptively. Intuitively, the importance distribution is concentrated in certain layers while relatively dispersed in other layers. As a suboptimal solution, retaining the same cache size for each layer results in obvious information loss in dispersed layers, and redundant cache in concentrated layers, as shown in Fig.˜1.(a).
Figure 1: Comparison between previous methods and ours. Previous methods often keep the same prefix length for priority sequences of KV, i.e., retraining the same cache size for each layer. This causes notable information loss for certain layers. In this example, the first layer loses 30% of information. In contrast, we derive the optimal global prefix configuration to preserve as much information as possible in each layer. In this example, both layers can retain 90% of information, thereby enhancing performance.
[…]
@@ 2 Related Work
2.1 Vision instruction-following model.
@@
The progress of large vision-language models (LVLMs) has significantly expanded the capabilities of large language models (LLMs) achiam2023gpt ; touvron2023llama ; touvron2023llama2 ; dubey2024llama ; jiang2023mistral ; bai2023qwen ; team2024gemma ; team2024gemma2 ; team2023gemini by incorporating visual information, leading to powerful generation and reasoning ability for multimodal tasks liu2024visual ; liu2024improved ; instructblip ; zhu2023minigpt ; dong2025fusion ; chen2023minigpt ; qiao2025lipt ; li2024mini ; beyer2024paligemma . These models typical use linear projection liu2024visual or perceivers jaegle2021perceiver to integrate visual representations into the input of LLMs directly. Then, they are further finetuned on high-quality instructional datasets, which include the image-text pairs and the language instructional commands. These models can thus follow multimodal instructions effectively and accurately. Inspired by the notable advancements, subsequent works have sought to enhance their abilities of region-level grounding  cai2023making ; zhang2023gpt4roi ; peng2023kosmos ; chen2023shikra , video comprehension zhang2023video ; lin2023video , industrial anomaly recognition gu2024anomalygpt ; yang2023dawn ; cao2023towards , and biomedical image understanding li2024llava ; van2024large , etc.
2.2 KV cache compression.
@@
While powerful, existing LVLMs typically encounter high inference costs due to the large parameter count and complex computation, hindering the practical applications. Given this, numerous efforts have focused on improving the inference efficiency, including quantization lin2024awq ; xiao2023smoothquant , distillation gu2024minillm ; shu2024llava , and pruning ma2023llm ; sreenivas2024llm ; huang2024dynamic , etc. These methods typically focus on the compression on the model level, reducing inference overhead brought by the parameters. Meanwhile, the KV cache on the data level also incurs significant memory usage and GPU communication overhead during inference. To address this, recent works have investigated various ways to compress the KV cache xiao2023efficient ; liu2024efficient ; he2024zipvl ; ge2023model ; ren2024efficacy ; wan2024look ; wang2024model . For example, StreamingLLM xiao2023efficient leverages the distance to the output token as the importance indicator and preserves starting KV vectors and those adjacent to the output token. H2zhang2024h2o utilizes the attention scores as importance metric for KV vectors and retains the important ones during generation. KVMerger wang2024model identifies sets of suitable KV states for merging and uses gaussian kernel weighted aggregation. Elastic Cache liu2024efficient divides the KV cache into several buckets according the positions of important KV cache and merges the vectors in the same bucket into one. While effective, they often overlook the distinct importance distribution across layers and retain the same KV cache size in each layer, causing notable information loss. Moreover, despite works like zhang2024pyramidkv ; ge2023model ; lee2024infinigen explore layer-wise KV cache size allocation, they often depend on manually designed patterns, which still achieve inferior model’s generation capability.
@@ 3 Methodology
In Sec.˜3.1, we first introduce the inference process of LVLMs and the general KV cache compression framework. Then, in Sec.˜3.2, we formalize the process of KV compression as retaining the prefix of KV cache and uncover the fact of diverse importance distributions of KV vectors in layers. In Sec.˜3.3, we further present PrefixKV, to deliver the ideal layer-wise cache size for the next token prediction by binary searching for the optimal global prefix configuration.
@@ 3.1 Preliminary
The general KV cache compression framework in existing works consist of two stages liu2024efficient ; zhang2024h2o ; xiao2023efficient   [^2] For more preliminary, please refer to previous works pope2023efficiently ; zhao2023survey ; yin2023survey ; zhou2024survey .. Firstly, after prefilling, the importance of each KV vector is derived based on the attention scores or the distance to the generated output. Then, the most important ones are retained in the cache while the less important KV vectors are removed to reduce the memory footprint. Meanwhile, the reserved KV cache sizes in layers meet the requirement of the compression ratio budget. Secondly, during decoding stage, with the inclusion of KV vectors in cache for newly generated tokens, the importance metric of each KV vector is updated. Less critical ones are removed to ensure that the cache size consistently aligns with the overall budget. Our method also follows this general framework, as shown in Fig.˜3. Specifically, after prefilling, we retain the most important KV vectors by the ideal global prefix configuration. In the decoding, we follow liu2024efficient to prune the vectors at a fixed distance to the latest generated token and maintain the global prefix configuration to meet target budget. Additionally, our method is also complementary to prefilling acceleration ways like FastV chen2024image .
@@ 3.2 Layer-wise KV Cache Importance Distribution
Mathematical notations for KV compression. We first introduce the necessary notations to formalize the compression procedure. Specifically, we follow liu2024efficient ; zhang2024h2o to employ the attention scores as the importance metric, which indicates the amount of contextual information of each KV vector. We suppose that the model consists of $L$ transformer layers. For the $l$-th, layer, its input tokens $\{\boldsymbol{t}_{l}^{1},\boldsymbol{t}_{l}^{2},...,\boldsymbol{t}_{l}^{N}\}$ interact with each other in the multi-head self-attention module, where $N$ is the token number. For the $i$-th head, we denote the query, key, and value vectors of the token $\boldsymbol{t}_{l}^{n}$ as $\boldsymbol{q}_{l}^{i,n}$, $\boldsymbol{k}_{l}^{i,n}$, and $\boldsymbol{v}_{l}^{i,n}$, respectively. Then, the causal attention score matrix $\boldsymbol{A}_{l}^{i}=\{\boldsymbol{a}_{l}^{i,m,n}\}_{N\times N}$ can be derived by $\boldsymbol{a}_{l}^{i,m,n}=\frac{\text{exp}(\boldsymbol{q}_{l}^{i,m}\cdot\boldsymbol{k}_{l}^{i,n})}{\sum_{j\leq m}\text{exp}(\boldsymbol{q}_{l}^{i,m}\cdot\boldsymbol{k}_{l}^{i,j})}$, where $\boldsymbol{a}_{l}^{i,m,n}$ indicates the attention score of token $\boldsymbol{t}_{l}^{m}$ with respect to token $\boldsymbol{t}_{l}^{n}$ in the $i$-th head. Then, the total attention score that each token $\boldsymbol{t}_{l}^{n}$ receives in the $i$-th head can be derived by $\sum_{m}\boldsymbol{a}_{l}^{i,m,n}$. For each $\boldsymbol{k}_{l}^{i,n}$ and $\boldsymbol{v}_{l}^{i,n}$, we follow liu2024efficient to define their importance metric as the averaged total attention score of $\boldsymbol{t}_{l}^{n}$ across all heads, i.e., $\boldsymbol{I}_{l}^{n}=\text{Average}_{i}(\sum_{m}\boldsymbol{a}_{l}^{i,m,n})$, where $\text{Average}_{i}$ denotes the average operation for heads. It is noted that the importance metric of each KV vector varies at each layer but is the same across heads. Then, for a compression ratio budget $r$, the top $\boldsymbol{R}_{l}$ proportion of KV vectors with the highest importance are retained in each head at the $l$-th layer. Besides, the adoption of $\{\boldsymbol{R}_{1},\boldsymbol{R}_{2},...,\boldsymbol{R}_{L}\}$ satisfy the requirement of $\sum_{l=1}^{L}\boldsymbol{R}_{l}N=rLN$.
@@ 3.3 PrefixKV
Based on above observations, we introduce PrefixKV. With layer-wise cache size scheme formalized as global prefix configuration, it employs binary search to derive ideal solution, as shown in Fig.˜3.(b).
Figure 3: (a) The inference process of LVLMs, where the orange and green rectangles denote the KV cache generated during prefilling and utilized during decoding, respectively. After prefilling, the KV cache is layer-wisely compressed according to the proportions specified by PrefixKV, i.e., $\{\boldsymbol{R}_{1},...,\boldsymbol{R}_{L}\}$. During decoding, as the sequence lengthens and cache increases, the KV cache consistently maintains the derived compression proportions by pruning KV at a fixed distance liu2024efficient . (b) The overview of PrefixKV. It employs binary search for cumulative priority sequences of KV to derive the optimal global prefix configuration, which delivers ideal cache size ratio for each layer.
Global prefix configuration. All layers’ cache size ratios $\{\boldsymbol{R}_{1},...,\boldsymbol{R}_{L}\}$, i.e., the prefix size ratios of KV cache, constitute the global prefix configuration space of the model. The goal is to identify the optimal global prefix configuration to maintain the high quality of model generation. Given the compression ratio budget $r$, we discover a information retention ratio $p$ to derive such configuration. Specifically, with the priority sequences, $p$ represents the cumulative priority threshold. For the $l$-th layer, its proportion of retained KV vectors is thus the minimum prefix size ratio $o$ such that the cumulative priority $\boldsymbol{P}_{l}^{o}$ is larger than or equal to $p$, i.e., $\boldsymbol{R}_{l}=\text{min}(\{o|\boldsymbol{P}_{l}^{o}\geq p\})$. Meanwhile, the value of $p$ satisfies the requirement of the compression ratio budget, i.e.,
@@
Initialize $p_{1}\leftarrow 0$, $p_{2}\leftarrow 1$;
while $p_{1}<p_{2}$ do
 $p\leftarrow\frac{p_{1}+p_{2}}{2}$, $\sum_{l}\boldsymbol{R}_{l}=\sum_{l}\text{min}(\{o|\boldsymbol{P}_{l}^{o}\geq p\})$;
 $\delta=\sum_{l}\boldsymbol{R}_{l}-rL$;
 if $\delta==0$ then return $p$ ;
 else if $\delta<0$ then $p_{1}\leftarrow p$ ;
 else $p_{2}\leftarrow p$ ;
return $p$;
Algorithm 1 Binary Search for ratio $p$
Binary search for optimal configuration. Due to the numerous possible values of $p$, obtaining $p$ is quite challenging. We propose employing binary search to efficiently derive $p$ for adaptive layer-wise KV retention. Formally, we start with the interval of $p_{1},p_{2}]$ where $p_{1}=0$ and $p_{2}=1$. We try $p=\frac{p_{1}+p_{2}}{2}=0.5$ and calculate its corresponding compression budget difference $\delta=\sum_{l}\boldsymbol{R}_{l}-rL$. If $\delta$ is equal to zero, we have found the value for $p$. If it is smaller than zero, we set $p_{1}=p$; otherwise, we set $p_{2}=p$. Then, we update $p=\frac{p_{1}+p_{2}}{2}$ and repeat the process until meeting the constraint. Algorithm˜1 presents the process. In this way, as shown in the dashed red line in Fig.˜2, different layers can consistently keep the maximal amount of the contextual information. It ensures that the valuable KV vectors are cached and accessible during generation across all layers. The quality of the model’s outputs can be maintained even after significant compression of the KV cache. In practice, $p$ that satisfies $\delta=0$ may not exist. We can thus set a small threshold for $\delta$ to terminate the search and scale the resulting global prefix configuration to meet the target budget.
Moreover, we observe that the cumulative priority sequences of layers are similar and robust across different samples. Therefore, given a compression ratio budget, we can leverage random samples to derive the corresponding $p$ value and the optimal global prefix configuration $\{\boldsymbol{R}_{1},\boldsymbol{R}_{2},...,\boldsymbol{R}_{L}\}$ for each layer offline, as analyzed in Fig.˜4, Tab.˜6, and Tab.˜5. The configuration can thus be adopted for the model during inference, which avoids the online binary search and shows good generalizability.